# Coherence length in a dilute flat-band superconductor

### v1

@@ I Introduction
In multiband superconductors, geometric effects, while inherent, are generally overshadowed by conventional effects in the BCS limit [1, 2]. However, as the bandwidth of a Bloch band narrows, conventional effects diminish due to their dependence on the derivative of the Bloch bands. In contrast, geometric effects, which are determined by the derivative of the Bloch states, can become dominant under these conditions. Inspired by recent technological advances in creating two-dimensional materials with nearly-flat electronic bands [16], there is a growing surge of interest in understanding geometric effects in multiband superconductors [17, 18, 19]. Among them, there has been a perplexing GL study regarding the size of Cooper pairs in a flat-band superconductor, suggesting they are limited by a fundamental length scale determined by the quantum geometry of the flat band [20]. However, subsequent work using the Bogoliubov-de Gennes approach for various lattice models in one and two dimensions has demonstrated that characteristic correlation lengths can be smaller than one lattice spacing without being constrained by the quantum metric [21].
To address these discrepancies [20, 21], we investigate the quantum-geometric effects on the GL coherence length in a dilute flat-band superconductor using a BCS-BEC crossover approach to the multiband pyrochlore-Hubbard model near the critical temperature for pairing. Our self-consistent results are in good agreement with the zero-temperature coherence length, demonstrating that GL coherence length decreases to zero monotonously as the interaction strength increases. In addition, we find that the effective mass of Cooper pairs aligns closely with that of the lowest-lying two-body bound states in the dilute regime. The pyrochlore lattice, akin to well-studied two-dimensional models like the Mielke-checkerboard and kagome lattices known for uniform pairing, provides an ideal framework for exploring three-dimensional flat-band superconductivity [6]. Recent demonstrations of flat bands and superconductivity in materials like pyrochlore metal CaNi2 [18] and pyrochlore superconductor CeRu2 [19] underscore the relevance of this model.
The rest of of the paper is organized as follows. In Sec. II, we introduce the multiband Hubbard Hamiltonian in reciprocal space, review the effective-action formalism near $T_{c}$, and derive the expansion coefficients for the time-dependent-GL (TDGL) theory. In Sec. III, we discuss self-consistency relations that are used to study dilute flat-band superconductivity in the pyrochlore lattice. In Sec. IV, we introduce the zero-temperature coherence length. In Sec. V, we present the numerical results and analyze them in various limits. The paper ends with a summary and outlook in Sec. VI, and the breakdown of zero-temperature coherence length is discussed in the Appendix.
@@ II.2 Effective action near $T\leq T_{c}$
In order to construct a microscopic TDGL theory near $T_{c}$ [22], we first extract the effective free-energy density of the system per lattice site as a function of the pairing order parameter from the effective action, i.e., from $\frac{T}{N}(S_{2}+S_{4}),$ where $T$ is the temperature in units of the Boltzmann constant $k_{\mathrm{B}}\to 1$, and $S_{2}$ ($S_{4}$) is the quadratic (quartic) contribution to the action [5]. The saddle-point action $S_{0}$ is not needed here because it does not depend on the order parameter at $T_{c}$, and $S_{1}$ and $S_{3}$ vanish at any $T$ due to the saddle-point condition.
@@ II.3 Coefficients for the TDGL expansion
Assuming this is the case, and by making an analogy with the Gross-Pitaevskii equation for a weakly-interacting atomic Bose gas, i.e., through a scaling of the order parameter [5, 22], we define $\mu_{p}(T)=a_{0}\epsilon(T)/d$ as the effective chemical potential of the Cooper pairs, $(M_{p}^{-1})_{ij}=c_{ij}/d$ as the inverse effective-mass tensor $\mathbf{M}_{p}^{-1}$ of the pairs, and $U_{pp}=b_{0}/d^{2}$ as the effective onsite repulsive interaction between the pairs. This identification suggests that $\mathbf{M}_{p}$ is composed of a conventional contribution $(M_{p}^{-1})^{\mathrm{intra}}_{ij}=c_{ij}^{\mathrm{intra}}/d$ and a geometric one $(M_{p}^{-1})^{\mathrm{inter}}_{ij}=c_{ij}^{\mathrm{inter}}/d.$ In this paper, we are primarily interested in the GL coherence length $(\xi_{T}^{2})_{ij}=(\xi_{\mathrm{GL}}^{2})_{ij}/\epsilon(T),$ whose temperature-independent prefactor is given by
@@
This relation again suggests that $\xi_{\mathrm{GL}}$ is composed of a conventional contribution $(\xi_{\mathrm{GL}}^{2})^{\mathrm{intra}}_{ij}=c_{ij}^{\mathrm{intra}}/(2a_{0})$ and a geometric one $(\xi_{\mathrm{GL}}^{2})^{\mathrm{inter}}_{ij}=c_{ij}^{\mathrm{inter}}/(2a_{0}).$ Since the calculation of $\xi_{\mathrm{GL}}$ requires $\mu$ and $T_{c}$ as inputs, we next describe our recipe for their self-consistent evaluation as a function of $U$, which is appropriate for a dilute flat-band superconductor in the pyrochlore lattice.
@@ III Self-consistency relations: $(\mu,T_{c})$
One can show quite generally that $\Omega_{0}=TS_{0}$ is the saddle-point and $\Omega_{2}=T\sum_{q}\ln\det[T\boldsymbol{\Gamma}^{-1}(q)]$ is the quadratic contribution. At $T_{c}$, while the former leads to the Fermi-Dirac (FD) distribution of an unbound (free) Fermi gas of $\uparrow$ and $\downarrow$ fermions, the latter contribution is typically split into $\mathcal{N}_{2}=\mathcal{N}_{bs}+\mathcal{N}_{sc}$  [22, 26]. While the bound-state contribution $\mathcal{N}_{bs}$ arises from the isolated poles of $\boldsymbol{\Gamma}(q)$, the continuum of two-particle excitations leads to the scattering contribution $\mathcal{N}_{sc}$ arising from the branch cut of the logarithm. Thus, $\mathcal{N}_{bs}=-T\sum_{q}\frac{\partial[\det\boldsymbol{\Gamma}^{-1}(q)]/\partial\mu}{\det\boldsymbol{\Gamma}^{-1}(q)}$ can be simplified considerably by noting that $\det\boldsymbol{\Gamma}^{-1}(q)\propto\Pi_{s}(\mathrm{i}\nu_{\ell}-\omega_{s\mathbf{q}}),$ where $\omega_{s\mathbf{q}}$ are the poles of $\boldsymbol{\Gamma}(q)$ determined by $\det\boldsymbol{\Gamma}^{-1}(\mathbf{q},\omega_{s\mathbf{q}})=0.$ For instance, since $\partial\omega_{s\mathbf{q}}/\partial\mu=-2$ in the BEC limit, one can approximate $\mathcal{N}_{bs}=-T\sum_{\ell s\mathbf{q}}2/(\mathrm{i}\nu_{\ell}-\omega_{s\mathbf{q}})=2\sum_{s\mathbf{q}}f_{\mathrm{BE}}(\omega_{s\mathbf{q}}),$ after summing over the Matsubara frequencies, where $f_{\mathrm{BE}}(x)=1/(e^{x/T}-1)=[\coth\big{(}\frac{x}{2T}\big{)}-1]/2$ is the Bose-Einstein (BE) distribution of a bound Bose gas of Cooper pairs. On the other hand, the branch-cut contribution is given by $\mathcal{N}_{sc}=\sum_{\mathbf{q}}\int_{-\infty}^{+\infty}\frac{d\omega}{\pi}f_{\mathrm{BE}}(\omega)\frac{\partial\delta(\mathbf{q},\omega)}{\partial\mu},$ where $\delta(\mathbf{q},\omega)=-\mathrm{Arg}[\det\boldsymbol{\Gamma}^{-1}(\mathbf{q},\omega+\mathrm{i}0^{+})]$ is the argument of the the propagator [22, 26]. For instance, if we decompose $\det\boldsymbol{\Gamma}^{-1}(\mathbf{q},\omega+\mathrm{i}0^{+})=R+\mathrm{i}I$ into its real and imaginary parts then $\frac{\partial\delta(\mathbf{q},\omega)}{\partial\mu}=(\dot{R}^{\mu}I-R\dot{I}^{\mu})/(R^{2}+I^{2}),$ where $\dot{R}^{\mu}=\partial R/\partial\mu$ and $\dot{I}^{\mu}=\partial I/\partial\mu.$ Since accurately computing $\mathcal{N}_{sc}$ is quite challenging, even in the context of much simpler continuum problems [22, 26], we next focus only on a dilute flat-band superconductor where $\mathcal{N}_{sc}$ is negligible for all $U\neq 0$, and can be omitted.
@@
In general, all of the two-body bound states can be determined from the isolated poles of Eq. (2) exactly, i.e., by computing $\det\boldsymbol{\Gamma}^{-1}(\mathbf{q},E_{s\mathbf{q}})=0,$ after setting the thermal factors $\mathcal{X}_{n\mathbf{k}\sigma}\to 1$ and $\mu\to 0$ [5]. However, in the presence of time-reversal symmetry and uniform pairing, the lowest-lying branch can simply be determined from the isolated pole of Eq. (3), i.e., by computing $\Gamma_{0}^{-1}(\mathbf{q},E_{1\mathbf{q}})=0,$ after again setting the thermal factors $\mathcal{X}_{n\mathbf{k}}\to 1$ and $\mu\to 0$. It can be shown that the overall structure of the two-body bound-state branches $E_{s\mathbf{q}}$ resemble the underlying Bloch bands $\varepsilon_{n\mathbf{k}}$ but with the opposite sign of energy [6]. For instance, $E_{1\mathbf{q}}$ resembles to $-\varepsilon_{1\mathbf{k}}$ but with an effective pair hopping parameter $t_{b}$ and some energy offset depending on $U$. To determine these effective parameters, we first expand $\varepsilon_{1\mathbf{k}}=-6\bar{t}+a^{2}\bar{t}|\mathbf{k}|^{2}/8$ in the low $\mathbf{k}\to\mathbf{0}$ limit, and identify the relation between the hopping parameter and the effective mass $m_{1}$ of the unpaired spin-$\sigma$ particles as $\bar{t}=4/(a^{2}m_{1})$. This suggests that $t_{b}=4/(a^{2}M_{b})$ is the relation between the effective hopping parameter and effective mass $M_{b}$ of the lowest-lying two-body bound states.
@@
In the low-$\mathbf{q}$ limit, we also recall that $E_{1\mathbf{q}}=E_{b}+\frac{1}{2}\sum_{ij}(M_{b}^{-1})_{ij}q_{i}q_{j}$  [15], where the energy offset $E_{b}$ is determined by $1=\frac{U}{N}\sum_{n\mathbf{k}}1/(2\varepsilon_{n\mathbf{k}}-E_{b}).$ Furthermore, very similar to that of the many-body bound states discussed in Sec. II.3, the inverse effective-mass tensor $(M_{b}^{-1})_{ij}=(M_{b}^{-1})^{\mathrm{intra}}_{ij}+(M_{b}^{-1})^{\mathrm{inter}}_{ij}$ of the lowest-lying two-body bound states is composed of an analogous conventional contribution $(M_{b}^{-1})^{\mathrm{intra}}_{ij}=\frac{1}{2D}\sum_{n\mathbf{k}}\ddot{\varepsilon}_{n\mathbf{k}}^{ij}/(2\varepsilon_{n\mathbf{k}}-E_{b})^{2}$ and an analogous geometric one $(M_{b}^{-1})^{\mathrm{inter}}_{ij}=\frac{1}{D}\sum_{n\mathbf{k}}g_{ij}^{n\mathbf{k}}/(2\varepsilon_{n\mathbf{k}}-E_{b})-\frac{1}{D}\sum_{n,m\neq n,\mathbf{k}}g_{ij}^{nm\mathbf{k}}/(\varepsilon_{n\mathbf{k}}+\varepsilon_{m\mathbf{k}}-E_{b}).$ It is important to emphasize that these two-body expressions are exact for any $U\neq 0$ in the pyrochlore lattice [6], and $(M_{b}^{-1})_{ij}=\delta_{ij}/M_{b}$ is diagonal as a consequence of uniform pairing. Thus, $E_{b}$ and $M_{b}$ can be determined self-consistently through these analytical expressions as functions of $U$, where $D=\sum_{n\mathbf{k}}1/(2\varepsilon_{n\mathbf{k}}-E_{b})^{2}.$ Once $E_{b}$ and $M_{b}$ are determined, we identify that

### v3

# Coherence length and quantum geometry in a dilute flat-band superconductor

@@ I Introduction
In multiband superconductors, geometric effects, while inherent, are generally overshadowed by conventional effects in the BCS limit [1, 2]. However, as the bandwidth of a Bloch band narrows, conventional effects diminish due to their dependence on the derivative of the Bloch bands. In contrast, geometric effects, which are determined by the derivative of the Bloch states, can become dominant under these conditions. Inspired by recent technological advances in creating two-dimensional materials with nearly-flat electronic bands [16], there is a growing surge of interest in understanding geometric effects in multiband superconductors [17, 18, 19]. Among them, there has been a perplexing zero-temperature GL study regarding the size of Cooper pairs in a flat-band superconductor, suggesting they are limited by a fundamental length scale determined by the quantum geometry of the flat band [20]. However, subsequent work using the Bogoliubov-de Gennes approach for various lattice models in one and two dimensions has demonstrated that characteristic correlation lengths can be smaller than one lattice spacing without being constrained by the quantum metric [21].
To address these discrepancies [20, 21], we investigate the quantum-geometric effects on the GL coherence length in a dilute flat-band superconductor using a BCS-BEC crossover approach to the multiband pyrochlore-Hubbard model near the critical temperature $T_{c}$ for pairing. Our self-consistent results are in good agreement with the zero-temperature coherence length, demonstrating that GL coherence length decreases to zero monotonously as the interaction strength increases. In addition, we find that the effective mass of Cooper pairs aligns closely with that of the lowest-lying two-body bound states in the dilute regime. The pyrochlore lattice, akin to well-studied two-dimensional models like the Mielke-checkerboard and kagome lattices known for uniform pairing, provides an ideal framework for exploring three-dimensional flat-band superconductivity [6]. Recent demonstrations of flat bands and superconductivity in materials like pyrochlore metal CaNi2 [18] and pyrochlore superconductor CeRu2 [19] underscore the relevance of this model.
The rest of of the paper is organized as follows. In Sec. II, we introduce the multiband Hubbard Hamiltonian in reciprocal space, review the effective-action formalism near $T_{c}$, and derive the expansion coefficients for the time-dependent GL (TDGL) theory. In Sec. III, we discuss self-consistency relations that are used to study dilute flat-band superconductivity in the pyrochlore lattice. In Sec. IV, we introduce the zero-temperature coherence length. In Sec. V, we present the numerical results and analyze them in various limits. In Sec. VI, we comment on the main source of discrepancy. The paper ends with a summary and outlook in Sec. VII, and the breakdown of zero-temperature coherence length is discussed in the Appendix.
@@ II.2 Effective action near $T\leq T_{c}$
In order to construct a microscopic TDGL theory near $T_{c}$ [22], we first extract the effective free-energy density of the system per lattice site as a function of the pairing order parameter from the effective action, i.e., from $\frac{T}{N}(\mathcal{S}_{2}+\mathcal{S}_{4}),$ where $T$ is the temperature in units of the Boltzmann constant $k_{\mathrm{B}}\to 1$, and $\mathcal{S}_{2}$ ($\mathcal{S}_{4}$) is the quadratic (quartic) contribution to the action [5]. The saddle-point action $\mathcal{S}_{0}$ is not needed here because it does not depend on the order parameter at $T_{c}$, and $\mathcal{S}_{1}$ and $\mathcal{S}_{3}$ vanish at any $T$ due to the saddle-point condition.
@@ II.3 Coefficients for the TDGL expansion
Assuming this is the case, and by making an analogy with the Gross-Pitaevskii equation for a weakly-interacting atomic Bose gas, i.e., through a scaling of the order parameter [5, 22], we define $\mu_{p}(T)=a_{0}\epsilon(T)/d$ as the effective chemical potential of the Cooper pairs, $(M_{p}^{-1})_{ij}=c_{ij}/d$ as the inverse effective-mass tensor $\mathbf{M}_{p}^{-1}$ of the pairs, and $U_{pp}=b_{0}/d^{2}$ as the effective onsite repulsive interaction between the pairs. This identification suggests that $\mathbf{M}_{p}$ is composed of a conventional contribution $(M_{p}^{-1})^{\mathrm{intra}}_{ij}=c_{ij}^{\mathrm{intra}}/d$ and a geometric contribution $(M_{p}^{-1})^{\mathrm{inter}}_{ij}=c_{ij}^{\mathrm{inter}}/d.$ In this paper, we are primarily interested in the GL coherence length $(\xi_{T}^{2})_{ij}=(\xi_{\mathrm{GL}}^{2})_{ij}/\epsilon(T),$ whose temperature-independent prefactor is given by
@@
This relation again suggests that $\xi_{\mathrm{GL}}$ is composed of a conventional contribution $(\xi_{\mathrm{GL}}^{2})^{\mathrm{intra}}_{ij}=c_{ij}^{\mathrm{intra}}/(2a_{0})$ and a geometric contribution $(\xi_{\mathrm{GL}}^{2})^{\mathrm{inter}}_{ij}=c_{ij}^{\mathrm{inter}}/(2a_{0}).$ Since the calculation of $\xi_{\mathrm{GL}}$ requires $\mu$ and $T_{c}$ as inputs, we next describe our recipe for their self-consistent evaluation as a function of $U$, which is appropriate for a dilute flat-band superconductor in the pyrochlore lattice.
@@ III Self-consistency relations: $(\mu,T_{c})$
One can show quite generally that $\Omega_{0}=T\mathcal{S}_{0}$ is the saddle-point and $\Omega_{2}=T\sum_{q}\ln\det[T\boldsymbol{\Gamma}^{-1}(q)]$ is the quadratic contribution. At $T_{c}$, while the former leads to the Fermi-Dirac (FD) distribution of an unbound (free) Fermi gas of $\uparrow$ and $\downarrow$ fermions, the latter contribution $\mathcal{N}_{2}=\sum_{\mathbf{q}}\int_{-\infty}^{+\infty}\frac{d\omega}{\pi}f_{\mathrm{BE}}(\omega)\frac{\partial\delta(\mathbf{q},\omega)}{\partial\mu},$ where $\delta(\mathbf{q},\omega)=-\mathrm{Arg}[\det\boldsymbol{\Gamma}^{-1}(\mathbf{q},\omega+\mathrm{i}0^{+})]$ is the argument of the propagator, is typically split into $\mathcal{N}_{2}=\mathcal{N}_{bs}+\mathcal{N}_{sc}$  [22, 26]. While the bound-state contribution $\mathcal{N}_{bs}$ arises from the isolated poles of $\boldsymbol{\Gamma}(q)$, the continuum of two-particle excitations leads to the scattering contribution $\mathcal{N}_{sc}$ arising from the branch cut of the logarithm. Suppose the branch point is at $\omega=\omega_{\mathbf{q}}^{*}$ below which $\delta(\mathbf{q},\omega)=0$. Thus, $\mathcal{N}_{bs}=-T\sum_{q}\frac{\partial[\det\boldsymbol{\Gamma}^{-1}(q)]/\partial\mu}{\det\boldsymbol{\Gamma}^{-1}(q)}$ can be simplified considerably by noting that $\det\boldsymbol{\Gamma}^{-1}(q)\propto\Pi_{s}(\mathrm{i}\nu_{\ell}-\omega_{s\mathbf{q}}),$ where $0\leq\omega_{s\mathbf{q}}<\omega_{\mathbf{q}}^{*}-2\mu$ are the poles of $\boldsymbol{\Gamma}(q)$ determined by $\det\boldsymbol{\Gamma}^{-1}(\mathbf{q},\omega_{s\mathbf{q}})=0.$ For instance, since $\partial\omega_{s\mathbf{q}}/\partial\mu=-2$ in the BEC limit where the bound state is made of two fermions with opposite spins, one can approximate $\mathcal{N}_{bs}=-T\sum_{\ell s\mathbf{q}}2/(\mathrm{i}\nu_{\ell}-\omega_{s\mathbf{q}})=2\sum_{s\mathbf{q}}f_{\mathrm{BE}}(\omega_{s\mathbf{q}}),$ after summing over the Matsubara frequencies, where $f_{\mathrm{BE}}(x)=1/(e^{x/T}-1)=[\coth\big{(}\frac{x}{2T}\big{)}-1]/2$ is the Bose-Einstein (BE) distribution of a bound Bose gas of Cooper pairs. Here, the $\mathbf{q}$ sum is restricted to only those center-of-mass momenta for which the corresponding bound states exist. This analysis is equivalent to approximating $\delta_{bs}(\mathbf{q},\omega)=\pi\sum_{s}\theta(\omega-\omega_{s\mathbf{q}})$ as the bound-state contribution to the phase shift near $\omega\approx\omega_{s\mathbf{q}},$ where $\theta(x)$ is the Heaviside step function [26, 22]. On the other hand, the branch-cut contribution is given by $\mathcal{N}_{sc}=\sum_{\mathbf{q}}\int_{\omega_{\mathbf{q}}^{*}}^{+\infty}\frac{d\omega}{\pi}f_{\mathrm{BE}}(\omega)\frac{\partial\delta(\mathbf{q},\omega)}{\partial\mu}.$ For instance, if we decompose $\det\boldsymbol{\Gamma}^{-1}(\mathbf{q},\omega+\mathrm{i}0^{+})=R+\mathrm{i}I$ into its real and imaginary parts then $\frac{\partial\delta(\mathbf{q},\omega)}{\partial\mu}=(\dot{R}^{\mu}I-R\dot{I}^{\mu})/(R^{2}+I^{2}),$ where $\dot{R}^{\mu}=\partial R/\partial\mu$ and $\dot{I}^{\mu}=\partial I/\partial\mu.$ Since accurately computing $\mathcal{N}_{sc}$ is quite challenging, even in the context of much simpler continuum problems [22, 26], we next focus only on a dilute flat-band superconductor where $\mathcal{N}_{sc}$ is negligible for all $U\neq 0$, and can be omitted.
@@
In general, all of the two-body bound states can be determined from the isolated poles of Eq. (2) exactly, i.e., by computing $\det\boldsymbol{\Gamma}^{-1}(\mathbf{q},E_{s\mathbf{q}})=0,$ after setting the thermal factors $\mathcal{X}_{n\mathbf{k}\sigma}\to 1$ and $\mu\to 0$ [5]. However, in the presence of time-reversal symmetry and uniform pairing, the lowest-lying branch can simply be determined from the isolated pole of Eq. (3), i.e., by computing $\Gamma_{0}^{-1}(\mathbf{q},E_{1\mathbf{q}})=0,$ after again setting the thermal factors $\mathcal{X}_{n\mathbf{k}}\to 1$ and $\mu\to 0$. It can be shown that the overall structure of the two-body bound-state branches $E_{s\mathbf{q}}$ resemble the underlying Bloch bands $\varepsilon_{n\mathbf{k}}$ but with the opposite sign of energy [6]. For instance, $E_{1\mathbf{q}}$ resembles to $-\varepsilon_{1\mathbf{k}}$ but with an effective pair hopping parameter $t_{b}$ and some energy offset depending on $U$. To determine these effective parameters, we first expand $\varepsilon_{1\mathbf{k}}=-6\bar{t}+a^{2}\bar{t}|\mathbf{k}|^{2}/8$ in the low $\mathbf{k}\to\mathbf{0}$ limit, and identify the relation between the hopping parameter and the effective mass $m_{1}$ of the unpaired spin-$\sigma$ particles as $\bar{t}=4/(a^{2}m_{1})$  111 We note that the matrix elements $(m_{n}^{-1})_{ij}=\ddot{\varepsilon}_{n\mathbf{k}_{0}}^{ij}$ of the inverse effective band-mass tensor $\mathbf{m}_{n}^{-1}$ of the $n$th Bloch band itself can also be written as $(m_{n}^{-1})_{ij}=(m_{n}^{-1})_{ij}^{\mathrm{intra}}+(m_{n}^{-1})_{ij}^{\mathrm{inter}},$ where $(m_{n}^{-1})_{ij}^{\mathrm{intra}}=\langle n_{\mathbf{k}_{0}}|\ddot{\mathbf{h}}_{\mathbf{k}_{0}}^{ij}|n_{\mathbf{k}_{0}}\rangle$ is the conventional intraband contribution and $(m_{n}^{-1})_{ij}^{\mathrm{inter}}=\sum_{m}(\varepsilon_{n\mathbf{k}_{0}}-\varepsilon_{m\mathbf{k}_{0}})g_{ij}^{nm\mathbf{k}_{0}}$ is the geometric interband contribution [32]. Here, $\mathbf{k}_{0}$ denotes the location of the band minimum or maximum of the corresponding Bloch band, and $\ddot{\mathbf{h}}_{\mathbf{k}_{0}}^{ij}=\frac{\partial^{2}\mathbf{h}_{\mathbf{k}}}{\partial k_{i}\partial k_{j}}\big{|}_{\mathbf{k}_{0}}$ This suggests that $t_{b}=4/(a^{2}M_{b})$ is the relation between the effective hopping parameter and effective mass $M_{b}$ of the lowest-lying two-body bound states.
@@
In the low-$\mathbf{q}$ limit, we also recall that $E_{1\mathbf{q}}=E_{b}+\frac{1}{2}\sum_{ij}(M_{b}^{-1})_{ij}q_{i}q_{j}$  [15], where the energy offset $E_{b}$ is determined by $1=\frac{U}{N}\sum_{n\mathbf{k}}1/(2\varepsilon_{n\mathbf{k}}-E_{b}).$ Furthermore, very similar to that of the many-body bound states discussed in Sec. II.3, the inverse effective-mass tensor $(M_{b}^{-1})_{ij}=(M_{b}^{-1})^{\mathrm{intra}}_{ij}+(M_{b}^{-1})^{\mathrm{inter}}_{ij}$ of the lowest-lying two-body bound states is composed of an analogous conventional contribution $(M_{b}^{-1})^{\mathrm{intra}}_{ij}=\frac{1}{2D}\sum_{n\mathbf{k}}\ddot{\varepsilon}_{n\mathbf{k}}^{ij}/(2\varepsilon_{n\mathbf{k}}-E_{b})^{2}$ and an analogous geometric one $(M_{b}^{-1})^{\mathrm{inter}}_{ij}=\frac{1}{D}\sum_{n\mathbf{k}}g_{ij}^{n\mathbf{k}}/(2\varepsilon_{n\mathbf{k}}-E_{b})-\frac{1}{D}\sum_{n,m\neq n,\mathbf{k}}g_{ij}^{nm\mathbf{k}}/(\varepsilon_{n\mathbf{k}}+\varepsilon_{m\mathbf{k}}-E_{b}).$ Using integration by parts, i.e., $\sum_{\mathbf{k}}\ddot{\varepsilon}_{n\mathbf{k}}^{ij}/(2\varepsilon_{n\mathbf{k}}-E_{b})^{2}=4\sum_{\mathbf{k}}\dot{\varepsilon}_{n\mathbf{k}}^{i}\dot{\varepsilon}_{n\mathbf{k}}^{j}/(2\varepsilon_{n\mathbf{k}}-E_{b})^{3},$ the former expression can alternatively be written in the form of Eq. (6). In addition, similar to $c_{ij}^{\mathrm{inter}}$, it is pleasing to note that $(M_{b}^{-1})^{\mathrm{inter}}_{ij}$ also does not have any contribution from the band touchings, i.e., the first sum cancels those touching contributions from the second sum whenever $\xi_{n\mathbf{k}}=\xi_{m\mathbf{k}}$ for any $n\neq m$. It is also worth emphasizing that these two-body expressions are exact for any $U\neq 0$ in the pyrochlore lattice [6], and $(M_{b}^{-1})_{ij}=\delta_{ij}/M_{b}$ is diagonal as a consequence of uniform pairing. Thus, $E_{b}$ and $M_{b}$ can be determined self-consistently through these analytical expressions as functions of $U$, where $D=\sum_{n\mathbf{k}}1/(2\varepsilon_{n\mathbf{k}}-E_{b})^{2}.$ Once $E_{b}$ and $M_{b}$ are determined, we identify that
@@
leading to $\varepsilon_{b\mathbf{q}}=6t_{b}-2t_{b}(1+\sqrt{1+\alpha_{\mathbf{q}}})$ as the effective dispersion, where $\alpha_{\mathbf{q}}$ is defined in Sec. II.1, and $\mu_{b}=2\mu-E_{b}\to 0^{-}$ as the effective chemical potential of pairs, such that $\varepsilon_{b\mathbf{q}}=|\mathbf{q}|^{2}/(2M_{b})$ reproduces the two-body physics exactly for all $U\neq 0$ in the low $\mathbf{q}\to\mathbf{0}$ limit. In Sec. V, we benchmark this approximate yet relatively complicated finite-temperature recipe against a much simpler zero-temperature one for the pyrochlore lattice, demonstrating very good agreement between the two.
IV Coherence length at $T=0$